OpenStudy (anonymous):

PLEASE HELP ME I WILL MEDAL AND FAN!!!!!!!!! A store is selling 5 notebooks for $15.00. At this rate, what is the cost of 8 notebooks? A. $15.05 B. $15.08 C. $15.13 D. $24.00

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Do your math.. It's D.

OpenStudy (anonymous):

15/5 to find the rate

OpenStudy (anonymous):

D for sure dude

OpenStudy (anonymous):

15/5

OpenStudy (anonymous):

then multiply by 8 and tada, your answer. :-)

OpenStudy (superflytyguy):

Each notebook is $3.00

OpenStudy (anonymous):

then that times 8

OpenStudy (anonymous):

funny everyone reply's to the easy stuff

OpenStudy (anonymous):

I have 1 more question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

shoot

OpenStudy (anonymous):

i dont see it

OpenStudy (anonymous):

A baseball pitcher allowed 6 runs in 9 innings. Which equation shows how to calculate the rate of runs per inning?

OpenStudy (anonymous):

easy

OpenStudy (bohotness):

do yuo have answer choices

OpenStudy (anonymous):

same

OpenStudy (anonymous):

9/6= x

OpenStudy (bohotness):

,,,,,,,,,,,,

OpenStudy (anonymous):

is that right?

OpenStudy (bohotness):

are you answering me ??

OpenStudy (bohotness):

i mena asking??

OpenStudy (bohotness):

i thinnk so if you asking me

OpenStudy (anonymous):

9/6=1.5 1.5*x=y

OpenStudy (anonymous):

i was asking him

OpenStudy (anonymous):

yeah here A. r=